OpenStudy (anonymous):

A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this

OpenStudy (anonymous):

I want to know for Part A if my answer is correct. My answer: 5x + 10y = 15 4x + 4y = 6

OpenStudy (anonymous):

Never mind. My answer is correct because my system of equations has the same solution just like the original one. Solution is (0, 1.5) which satisfies both equations

OpenStudy (campbell_st):

I'd multiply the 2nd equation by -4 -8x -8y = -12 then the 2 equations 3x + 8y = 12 -8x - 8y = -12 --------------- -5x = 0 this is the sum of the 2 equations... it eliminated y that's my best guess

OpenStudy (anonymous):

How about mine is that okay?

OpenStudy (anonymous):

Only confusing part are the instructions. "replacing one equation with the sum of that equation and a multiple of the other."

OpenStudy (campbell_st):

your equations are equivalent but don't really help in obtaining a solution.

OpenStudy (campbell_st):

the fact that you were asked to sum the equations... is normally what is done to eliminate a variable